OpenStudy (anonymous):

(09.03) Nanna's Italian Restaurant has collected data about customer sauce orders. It calculated that P(marinara) = 0.64, P(alfredo) = 0.56, and P(marinara or alfredo) = 0.89. Determine the P(marinara and alfredo). 0.33 0.60 0.31 0.68

OpenStudy (anonymous):

@kirbykirby @satellite73 Can anyone help walk me through this one?

OpenStudy (anonymous):

\[P(A\cup B)=P(A)+P(B)-P(A\cap B)\]

OpenStudy (kirbykirby):

So let's abbreviate, M = marinara, A = alfredo What they tell you is: \(P(M) = 0.64\\ P(A)= 0.56\\ P(M \text{ or }A)=P(M \cup A)=0.89\) They want \(P(M \text{ and } A)=P(M \cap A)\) Use the rule: \(P(M \cup A ) = P(M) + P(A) - P(M \cap A)\)

OpenStudy (anonymous):

Ok give me just a minute

OpenStudy (anonymous):

or in @kirbykirby notation \[P(M\cap A)=P(M)+P(A)-P(M\cup A)\] when you solve for what you are looking for

OpenStudy (anonymous):

You are solving for p(mUa). ?

OpenStudy (anonymous):

Upside down u I mean (intersection)

OpenStudy (kirbykirby):

yes

OpenStudy (anonymous):

\cup is \(\cup\) union \cap is \(\cap\) intersection if you want to write it

OpenStudy (anonymous):

Ok so it is p(MUA)=P(0.64)+P(0.56)-P(0.89)

OpenStudy (anonymous):

So P(MUA)=0.31 @kirbykirby

OpenStudy (kirbykirby):

yup !

OpenStudy (anonymous):

Thank you